OpenStudy (anonymous):

solve the following inequality: (x^2(5+x)(x-5))/((x+2)(x-2)) >or equal to 0

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

I'm sorry, i probably can't do it. @freethinker

OpenStudy (anonymous):

\[\left( \frac{ x ^{2}(5+x)(x-5) }{ (x+2)(x-2) }\ge0 \right)\]

OpenStudy (freethinker):

can you simplify?

OpenStudy (anonymous):

Honestly, I wouldnt know where to start to simplify this

OpenStudy (freethinker):

maybe because it cannot be simplified? lol

OpenStudy (anonymous):

Good answer lol

OpenStudy (freethinker):

where are the math gods when you need them?

OpenStudy (freethinker):

or goddess *

OpenStudy (anonymous):

=/ yeah

OpenStudy (anonymous):

i think it can be done u just need to test a lot of intervals

OpenStudy (freethinker):

how about you just multiply both sides by x^2-4 to get rid of the fractions?

OpenStudy (anonymous):

intervals are determined by the zeros of the numerator and the zeros of the denominator

OpenStudy (freethinker):

good thinking julian

OpenStudy (anonymous):

x = 2, x = -2 for the denominator?

OpenStudy (anonymous):

Ok so wat you need to is.. divide and multiply by (x+2) and (x-2) .. so as to make the denominator (x+2)^2 (x-2)^2 .. that is +ve always.. so you can equate it with 0.. and the numerator will be x^2(5+x)*(x-5)(x+2)(x-2).. also take x^2 to the other side.. and you'll be left with.. (5+x)(x-5)(x+2)(x-2)... get it till now?

OpenStudy (anonymous):

(5+x)(x-5)(x+2)(x-2) >=0 *

OpenStudy (anonymous):

you cant multiply with x^2-4 because you dont know if its =ve or -ve.. for x>2.. it will be -ve and the inequality will be reversed.

OpenStudy (anonymous):

+ve*

OpenStudy (anonymous):

Rnr is doing well i think that freethinker r losing some intervals

OpenStudy (anonymous):

@musicalrose get what i did there?

OpenStudy (anonymous):

yes just not sure what to do with it, to me it looks like it cancels out to being no solution

OpenStudy (anonymous):

yup. ^

OpenStudy (anonymous):

from here (5+x)(x-5)(x+2)(x-2) >=0 we just need to test the intervals (-INF, -5)(-5,-2)(-2,2)(2,5)(5, INF)

OpenStudy (anonymous):

@freethinker its ok :) this problem just sucks

OpenStudy (anonymous):

so, if it is no solution how do I say that? my problem only asks "what is the solution" with an answer box, so I worry if there wasn't one, I would have a multiple choice.

OpenStudy (anonymous):

@julian25 i think it should be (-inf,-5][-2,2][5,inf)... for [-5,-2] and [2,5].. the expression would be negative..

OpenStudy (raden):